OpenStudy (anonymous):

A diver drops from a 10 meter high cliff. At what speed does he enter the water, and how long is he in the air?

OpenStudy (ankit042):

just use s=ut+1/2at^2 u=0;a=10;s=10

OpenStudy (anonymous):

use equation of continuity \[\rho gh+1/2\rho v ^{2}+P=const.\] rho=density of water P=pressure ,v = speed h =relative height

OpenStudy (ybarrap):

I like using potential energy = kinetic energy \[mgh=\frac{ 1 }{ 2 }mv ^{2}\] \[gh=\frac{ 1 }{ 2 }v ^{2}\] \[v ^{2}=2gh\] So speed = \[v=\sqrt{2gh}\] Using speed to calculate time in the air: \[h=vt\] \[t=\frac{ h }{ v }\]

OpenStudy (anonymous):

if the diver drops, shouldn't the diver be following the typical accelaration of y=yi+vi*t+1/2at^2. Since he drops, he accelerates at 9.81 m/s^2 and starts at a velocity of 0. Therefore y=10-1/2*a*t^2. Since y=10, 10=1/2*9.81*t^2 so (20/9.81)=t^2 and sqrt(20/9.81)=t. The overall algebraic equation should then be sqrt((2*y)/a)=t considering that the diver starts from rest and a=-9.81m/s^2

OpenStudy (anonymous):

See above: a=-9.81 m/s^2 v(initial)=0; y(initial)=10 Overall equation \[y=yi+vi*t+1/2*a*t ^{2}\] Velocity Equation \[ \sqrt{(2*y)/a}=t \]